# Time Series Foundation Models: Benchmarking Challenges and Requirements

### v1

@@ Abstract
Time Series Foundation Models (TSFMs) represent a new paradigm for time series forecasting, offering zero-shot forecasting capabilities without the need for domain-specific pre-training or fine-tuning. However, as with Large Language Models (LLMs), evaluating TSFMs is tricky, as with ever more extensive training sets, it becomes more and more challenging to ensure the integrity of benchmarking data. Our investigation of existing TSFM evaluation highlights multiple challenges, ranging from the representativeness of the benchmark datasets, over the lack of spatiotemporal evaluation, to risks of information leakage due to overlapping and obscure datasets, and the memorization of global patterns caused by external shocks like economic crises or pandemics. Our findings reveal widespread confusion regarding data partitions, risking inflated performance estimates and incorrect transfer of global knowledge to local time series. We argue for the development of robust evaluation methodologies to prevent pitfalls already observed in LLM and classical time series benchmarking, and call upon the research community to design new, principled approaches, such as evaluations on truly out-of-sample future data, to safeguard the integrity of TSFM assessment.
@@ 1 Introduction
Time Series Foundation Models (TSFMs) are an emerging class of time series forecasting models inspired by the architecture and training procedures of foundation models in natural language processing, especially Large Language Models (LLMs). Prominent examples of TSFMs include Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024) and Time-MoE (Shi et al., 2024). In contrast to traditional time series forecasting models, which need to be trained from scratch on the time series that should be forecasted, TSFMs use transfer learning. That is, they are pre-trained on massive amounts of generic and/or domain-specific time series and are, thus, able to produce zero-shot forecasts.
However, pre-training TSFMs on massive amounts of time series data, typically scraped from publicly accessible online repositories, introduces challenges similar to those observed in LLM evaluation. The practice of training LLMs on much of the Internet has led to an LLM “evaluation crisis" (Liao & Xiao, 2023), describing a situation in which most benchmark data has already been exposed to an LLM during pre-training. Combining such test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) with memorization effects (Chang et al., 2024) can result in overly-optimistic estimates of a model’s predictive performance and complicate the fair comparison of models through common benchmarks. Furthermore, as with LLMs (Chang et al., 2024), cross-validation in TSFMs is omitted due to its exhaustive resource requirements with large foundation models (Das et al., 2024; Gruver et al., 2024), and the diversity and generalizability of benchmarking datasets can be questioned (Aksu et al., 2024).
As the awareness about evaluation challenges in TSFMs is growing, researchers have begun to propose mitigation strategies, such as held-out-driven benchmarks (Aksu et al., 2024) and clean train-test splits within benchmarks (Qiu et al., 2024) that include more diverse and generalizing test sets Aksu et al. (2024). Especially when comparing TSFMs against each other, current approaches face limitations in defining an information-leakage-free test set for all investigated models (Aksu et al., 2024). This challenge is even more pronounced for larger models developed based on scaling laws, as their increased capacity may amplify the memorization effect, leading to a greater impact of test set contamination (see Section 2).
In this article, we discuss the evaluation challenges arising from current practices of training and evaluating TSFMs and propose requirements that a TSFM benchmark should fulfill. Empirically, our position is supported by an analysis of the lineage of the training and test sets of 15 prominent TSFMs and a further investigation of existing TSFM benchmarks [^1] TSFM benchmarks that existed in the beginning of 2025.. In addition, we provide a theoretical derivation of the phenomenon of global pattern memorization, which is specific to TSFMs and, to the best of our knowledge, has not been formally introduced before.
First, while we did not identify cases of information leakage within any of the underlying studies, our analysis highlights the increasing confusion with regards to a clear separation between training and test sets, one of the fundamental principles of time series forecasting and machine learning, and outlines possible adverse effects. Concretely, we found that one model’s training set is another model’s test set. This confusion, which is further exacerbated through the practice of remixing and renaming of datasets, stems from an ever-growing size of datasets, combined with non-transparent documentation about which exact data points have been used for training and testing, and complicates the fair benchmarking between TSFMs. We will show that this is not just a danger, but is actually already happening, using two examples where information leakage occurs in benchmarks 3.1.
Second, we highlight a more subtle instance of information leakage, particularly in time series forecasting, namely the memorization of global patterns. Global external shocks like the COVID-19 pandemic or global economic crises can profoundly impact numerous aspects of our economies and societies and, hence, are reflected in many diverse time series. A TSFM trained and evaluated on time series that are disjunct but stem from the same period and are both impacted by such an external shock can still profit from memorizing these patterns.
In addition to the issue of information leakage, we point out that current TSFM evaluation methods are not fully consistent with previous approaches in time series forecasting. In Section 3.3 and 3.4 we show that TSFMs typically do not implement best practices such as time series cross-validation or choosing representative time series for a domain.
This article argues that in the age of TSFMs a new benchmarking methodology is required to prevent a repetition of the LLM evaluation crisis and calls for new proposals that fulfill the requirements developed in this paper.
2 Foundation Models in Time Series
@@
A foundation model is “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks" (Bommasani et al., 2021). The term is primarily associated with Transformer-based LLMs, especially decoder-only models like the ChatGPT or Llama families of models (OpenAI, 2023; Grattafiori et al., 2024). The success of foundation models can be attributed to scaling laws (Kaplan et al., 2020), that is, increasing training data, model parameters, and computational resources significantly improves the model’s performance.
LLMs have demonstrated remarkable zero-shot learning capabilities across diverse benchmarks (Liang et al., 2022). However, concerns have been raised that these models might just memorize publicly available questions and answers of well-known benchmarks (Chang et al., 2024). Memorization (Carlini et al., 2022) is closely related to test set contamination, describing the situation where exact or highly similar samples from a test set are leaked to the LLM’s pre-training data (Li et al., 2024a). This contamination is a side-effect of the vast size and incomprehensible nature of today’s pre-training datasets, such as Common Crawl [^2] https://commoncrawl.org/ or The Pile [^3] https://pile.eleuther.ai/ (Li et al., 2024a; Singh et al., 2024). Despite efforts to improve dataset curation in recent LLMs (Grattafiori et al., 2024), evidence shows that test set contamination remains prevalent in various forms inflating generalization performance (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a).
Like generative LLMs, time series forecasting models are trained to autoregressively predict the next value(s) given a sequential series of past data points. Hence, it is not surprising, that researchers started to adapt decoder-only Transformers for time series forecasting. Some of these models encoded numerical time series values as natural language tokens (Gruver et al., 2024; Xue & Salim, 2024), others have developed novel tokenization techniques for numerical data, such as TimeGPT-1 (Garza et al., 2024) and Chronos (Ansari et al., 2024). Like LLMs, decoder-only Transformers for time series forecasting can be pre-trained on generic time series and adapted or fine-tuned to domain-specific time series optionally. At this, they seem to leverage the same scaling laws as LLMs (Edwards et al., 2024), leading to Time Series Foundation Models able to zero or few-shot forecast time series of different frequencies and domains adapting to diverse lengths of input context and forecasting horizons (Das et al., 2024; Liang et al., 2024).
@@
Traditional statistical or machine-learning-based time series forecasting models have to be trained on the time series that should be forecasted. These local models exhibit only limited generalization capabilities, that is, they can only predict values that are independent and identically distributed (i.i.d.) to their training data. Combined with the common practice to use time-based train/test splits when evaluating time series models (Hyndman & Athanasopoulos, 2021), this property of local models protects them from information leakage.
Table 1: Summary of models, citations, and release years.
Model Name Author Year
Time-MoE Shi et al. (2024) 2025
Chronos Ansari et al. (2024) 2024
Lag-LLama Rasul et al. (2024) 2024
Moirai Woo et al. (2024) 2024
Moirai-MoE Liu et al. (2024b) 2024
MOMENT Goswami et al. (2024) 2024
Timer Liu et al. (2024c) 2024
TimesFM Das et al. (2024) 2024
Time-LLM Jin et al. (2024) 2024
TTM Ekambaram et al. (2024) 2024
UniTime Liu et al. (2024a) 2024
VisionTS Chen et al. (2024) 2024
ForecastPFN Dooley et al. (2023) 2023
GPT4TS Zhou et al. (2023) 2023
LLMTime Gruver et al. (2024) 2023
TimeGPT-1 Garza et al. (2024) 2023
TSFMs, in contrast, are often evaluated on zero-shot or few-shot forecasting tasks using held-out datasets, similar to LLM evaluation (Ansari et al., 2024; Liu et al., 2024b; Woo et al., 2024). Notably, the terms “zero-shot” and analogously “few-shot” need to be distinguished between LLMs and TSFMs: while LLMs adapt to different tasks without any example in zero-shot learning (Brown et al., 2020), TSFMs usually forecast zero-shot by adapting to different input lengths, frequencies, horizons and domains (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024).
The specialized practices of pre-training and zero-shot forecasting in TSFMs introduce novel challenges while simultaneously amplifying existing difficulties in the evaluation of time series models, as discussed in the following sections.
@@ 3 Challenges of Evaluating Time Series Foundation Models
3.1 Challenge: Test Set Contamination Through Multi-Purpose Use of Datasets
In order to empirically study whether test set contamination is a problem in existing TSFM evaluations, we tried to track the lineage of the various datasets used for training and evaluating the TSFM listed in Table 1. Specifically, we distinguish between three types of data usages: pre-training datasets, train/test datasets, and zero-shot datasets.
Pre-training datasets are utilized exclusively for foundational training to teach typical time series patterns (e.g., trends, seasonality) to a model. Here, the entire time series is usually used for training, and no evaluation is conducted on these datasets. Train/test datasets, also called in-domain or in-distribution datasets, fit a model to time series from a specific domain (e.g., finance, energy, transportation). Following the traditional time series forecasting methodology, these datasets are typically split at a particular time point into training and test sets (Hyndman & Athanasopoulos, 2021; Kapoor & Narayanan, 2023). The training data is included in the model’s training corpus during pre-training or subsequent task-specific adaptation/fine-tuning. The test data is used to evaluate the model’s performance within the given domain. When benchmarking multiple models in a specific domain, all models should use the same datasets with the same train-test splits in order to ensure a fair comparison. Zero-shot datasets (held-out datasets) are entirely excluded from the pre-training and task-specific fine-tuning process and used for inference only. Unlike LLM zero-shot datasets with diverse tasks, TSFM datasets often share frequencies and domains with pre-training data, primarily evaluating frequency and domain generalization. However, in zero-shot evaluation, the TSFMs make forecasts, typically in a rolling manner, taking a limited context of historical values as inputs and producing output for future time steps (Aksu et al., 2024). Ideally, all TSFMs would be benchmarked on the same zero-shot datasets.

### v2

# Challenges and Requirements for Benchmarking Time Series Foundation Models

@@ Abstract
Time Series Foundation Models (TSFMs) represent a new paradigm for time-series forecasting, promising zero-shot predictions without the need for task-specific training or fine-tuning. However, similar to Large Language Models (LLMs), the evaluation of TSFMs is challenging: as training corpora grow increasingly large, it becomes difficult to ensure the integrity of the test sets used for benchmarking. Our investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series. Ignoring these forms of information leakage when benchmarking TSFMs risks producing overly optimistic performance estimates that fail to generalize to real-world settings. We therefore argue for the development of novel evaluation methodologies that avoid pitfalls already observed in both LLM and classical time-series benchmarking, and we call on the research community to adopt principled approaches to safeguard the integrity of TSFM evaluation.
@@ 1 Introduction
Time Series Foundation Models (TSFMs) represent an emerging paradigm in forecasting, drawing inspiration from the architecture and training methodologies of foundation models in natural language processing (NLP). Despite being a relatively new research area, the field has evolved into a highly dynamic landscape featuring a wide variety of architectures and approaches. This rapid proliferation is exemplified by diverse models such as Chronos (Ansari et al., 2024a), Chronos-bolt (Ansari et al., 2024b), Chronos-2 (Ansari et al., 2025), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024), Time-MoE (Shi et al., 2024), TimesFM (Das et al., 2024), TimesFM-2.0 (Das et al., 2024), TimesFM-2.5 (Das et al., 2024), or TiRex (Auer et al., 2025). In contrast to traditional time series models that are trained from scratch for specific tasks and time series, TSFMs leverage transfer learning: they are pre-trained on large collections of generic time series and can perform zero-shot forecasting without domain adaptation or fine-tuning (Liang et al., 2024).
The reliance on massive amounts of time series training data, typically scraped from public repositories, raises the question of whether the proper evaluation of TSFMs faces similar challenges to those observed in the evaluation of Large Language Models (LLMs). In the NLP domain, training on vast portions of the internet has given rise to an “evaluation crisis" (Liao and Xiao, 2023), in which test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) and memorization effects (Chang et al., 2024) have led to overly optimistic performance estimates. Whether current TSFM evaluations are vulnerable to analogous issues remains an open question, with important implications for fair and reliable model comparison.
While benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, we doubt that these approaches effectively prevent information leakage in the context of TSFMs. Beyond the risk of direct information leakage, the increased memorization capacity of scaled foundation models poses additional challenges (see Section 3).
This article examines the validity and reliability of current TSFM evaluation practices and discusses the risk of an evaluation crisis in the TSFM domain. We address this issue through two contributions: first, we conduct a comprehensive train-test set lineage analysis of 22 leading TSFMs and review all major evaluation benchmarks available as of early 2026. Second, we empirically demonstrate the phenomenon of temporal overlaps of correlated train and test series, a challenge specific to TSFMs, through a dedicated experiment.
Our investigation yields two primary findings that suggest that TSFM evaluation is indeed compromised:
  • First, while we find no evidence of unintended direct information leakage within the training and evaluation of individual TSFMs, our analysis reveals substantial ambiguity in data separation between models. In several cases, one model’s training set serves as another’s test set. This issue is further exacerbated by dataset remixing, renaming, and inadequate documentation, which already pose significant evaluation challenges for newer models (see Section 3.1).
  • Second, we identify the temporal overlap of correlated train and test series as an indirect form of information leakage specific to TSFMs. External shocks, such as the COVID-19 pandemic to stock prices, as well as exogenous drivers like local weather conditions to local solar production, act as factors that directly shape or distort the causal dynamics between multiple time series (see Section 3.2). As a result, a TSFM trained and evaluated on disjoint time series covering the same time period can effectively memorize patterns induced by such causal relations (see Figure 3).
Based on these findings, we argue that the rise of TSFMs necessitates new benchmarking methodologies to prevent an evaluation crisis akin to that observed in NLP.
2 Background
2.1 Time Series Foundation Models
@@
A foundation model is defined as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted or fine-tuned to a wide range of downstream tasks” (Bommasani et al., 2021). The term is primarily associated with Transformer-based Large Language Models (LLMs) like the ChatGPT or Llama families (OpenAI, 2023; Grattafiori et al., 2024). Their success is largely attributed to scaling laws (Kaplan et al., 2020), describing the practice of increasing training data, parameters, and compute to improve a model’s predictive performance. Although LLMs demonstrate impressive zero-shot capabilities across benchmarks (Liang et al., 2022), concerns persist that they partly rely on memorizing benchmark data (Chang et al., 2024; Carlini et al., 2022). Such test set contamination, where test samples appear in pre-training data like Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/, remains common despite improved dataset curation, leading to inflated performance estimates (Li et al., 2024a; Singh et al., 2024; Grattafiori et al., 2024; Mirzadeh et al., 2024; Ravaut et al., 2024).
Recognizing the structural similarity between language modeling and time series analysis, recent research has adapted the Transformer architecture to forecast future values of time series, ranging from the classical encoder-decoder Transformer (Ansari et al., 2024a) over Decoder-only models (Das et al., 2024) to Encoder-only models often applied in multivariate time series forecasting (Goswami et al., 2024). Beyond this, TSFMs are based on a variety of other architectural backbones such as LLMs Gruver et al. (2024); Xue and Salim (2024), Vision-based Autoencoders Chen et al. (2024), and mixed architectures that leverage LLM reprogramming Jin et al. (2024). Across these approaches, a key commonality lies in how data are represented: LLM-based models encode numerical values directly as language tokens like LLMTime (Gruver et al., 2024) and PromptCast (Xue and Salim, 2024), whereas some Transformer-based variants rely on specialized tokenization schemes to discretize continuous time series, as exemplified by TimeGPT-1 (Garza et al., 2024) and the T5-based Chronos (Ansari et al., 2024a). Furthermore, Vision-based approaches represent time series as gray-scaled images, while reprogramming-based approaches mix textual and continuous time series representations Jin et al. (2024). Similar to LLMs, all presented architectures are pre-trained on large collections of generic time series and appear to follow scaling laws (Edwards et al., 2024), enabling them to produce zero-shot forecasts and handle a wide range of frequencies and forecasting horizons. Only one exception—MOMENT—requires task-adaptive fine-tuning for high-quality forecasts Goswami et al. (2024). Based on an encoder-only architecture this model is capable of time series regression, classification, anomaly detection, and forecasting Goswami et al. (2024). However, due to its broad application area, the model needs task-adaptive fine-tuning of the last neural network layer that is specialized for each task. A key distinction between the foundation models in NLP and time series analysis concerns the definition of “zero-shot” capabilities. While LLMs typically adapt to entirely different tasks (Brown et al., 2020), TSFMs typically perform zero-shot forecasting by generalizing across domains, time series, input lengths, frequencies, and horizons without the need for model retraining (Ansari et al., 2024a; Das et al., 2024; Woo et al., 2024).
2.2 Benchmarking Methodologies
@@
Traditional statistical or machine-learning-based models for time series forecasting generally require explicit training on the target time series. These local models exhibit only limited generalization capabilities; that is, they can only predict values that are identically distributed to their training data. Consequently, the dominant logic in time series benchmarking relies on time-based train/test splits (Hyndman and Athanasopoulos, 2021). This approach creates a temporal separation between training and evaluation data to prevent information leakage and is employed in major time series collections such as the Monash repository (Godahewa et al., 2021) and TFB (Qiu et al., 2024).
To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice. Instead of using a single time-based split, TSCV uses the idea of rolling windows (see left in Figure 1). This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Makridakis et al., 2022; Bojer and Meldgaard, 2021). These competitions are typically held at irregular intervals.
While all of the above evaluation strategies rely on historical data for testing, the recent M6 competition introduced a live evaluation using future values of financial assets on selected dates as the target (Makridakis et al., 2024).
Figure 1: Evaluation strategies along the time (left) or the domain dimension (right).
While current TSFM benchmarking studies also rely on historical evaluation data, their logic diverges from the above described time-series cross-validation strategy. Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024). Figure 1 (b) shows the focus of most TSFM evaluations. Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024a). This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets. Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al. (2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks. Roque et al. (2025) showed that with just four cherry-picked test sets, out of the distribution of available test sets, 46% of models could appear as state-of-the-art purely due to selection bias. Notably, deep learning models were found to be more susceptible to this variance than classical baselines. Especially when evaluation general-purpose TSFMs, generalization is a core evaluation focus.
Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020). While this approach has first adopters, notably Goktas et al. (2025b); Shchur et al. (2025), it is not yet commonly used in TSFM benchmarking.
The scaling of pre-training and evaluation datasets in TSFMs gives rise to another challenge that can also be observed in LLM evaluation. As LLMs are pre-trained on large datasets crawled from the internet like Common Crawl [^3] https://commoncrawl.org/ or The Pile [^4] https://pile.eleuther.ai/, they have already seen most benchmark tasks during pre-training. Ultimately, this led to an LLM “evaluation crisis” Liao and Xiao (2023). Likewise, with large and obscure pre-training datasets, defining test datasets that do not overlap with models’ pre-training datasets becomes increasingly difficult in TSFMs Aksu et al. (2024); Goktas et al. (2025a); Xu et al. (2025). As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance of TSFMs Aksu et al. (2024). Thus, investigating information leakage in TSFMs is an important research direction. Addressing this, current benchmarks developed strategies to prevent information leakage. GIFT-Eval provides an explicit pretraining dataset containing approximately 230 billion data points and a dedicated test with 177 million data points Aksu et al. (2024) staying with a series cross validation (see Figure 1 (a)) but no time cross validation. Two other benchmarks TempusBench and Fidel-TS, use test datasets from unknown or restricted sources to prevent information leakage Goktas et al. (2025a); Xu et al. (2025). Particularly, TempusBench leverages a new test set with 48 uni- and multi-variate time series forecasting tasks that is currently not in any TSFM pre-training corpus Goktas et al. (2025a). While they apply a time and series cross-validation strategy, they further propose to leverage real-world live data for evaluations Goktas et al. (2025a). Fidel-TS apply a similar strategy by leveraging restricted live APIs but implement a series cross validation without time cross validation Xu et al. (2025).
Overall, all current approaches to mitigate information leakage and the possibly resulting inflation of performance rely on the development of new datasets that can currently not be in the pre-training corpus of TSFMs.
@@ 3 Challenges of Evaluating Time Series Foundation Models
Based on the need for an examination of information leakage in current TSFMs, we decided to first conduct a TSFM train-test dataset lineage analysis (Section 3.1). We then further analyze the outcomes of potential indirect overlaps in time series between pre-training and test sets of TSFMs.
3.1 Direct Information Leakage: Train-Test Sample Overlap
To examine potential test set contamination in existing TSFM evaluations, we searched for peer-reviewed and accepted papers including the terms "Time Series Foundation Model" at leading venues in the machine learning area (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops, published up to January 2026. This search resulted in 22 published TSFMs that met our inclusion criteria (Table 1). For each of these papers, we analyzed their dataset lineage, categorizing the used datasets into pre-training, train/test and zero-shot.
Table 1: Time series foundation models published in recent years
Model Name Author Year Conference/Journal
ForecastPFN Dooley et al. (2023) 2023 NeurIPS 23
GPT4TS Zhou et al. (2023) 2023 NeurIPS 23
Lag-Llama Rasul et al. (2024) 2024 NeurIPS 23 Workshop R0-FoMo
LLMTime Gruver et al. (2024) 2024 NeurIPS 23
Time-LLM Jin et al. (2024) 2024 ICLR 24
Chronos Ansari et al. (2024a) 2024 TMLR
TimesFM Das et al. (2024) 2024 ICML 24
Moirai Woo et al. (2024) 2024 ICML 24
MOMENT Goswami et al. (2024) 2024 ICML 24
Timer Liu et al. (2024c) 2024 ICML 24
UniTime Liu et al. (2024a) 2024 WWW 24
TinyTimeMixer Ekambaram et al. (2024) 2024 NeurIPS 24
Time-MoE Shi et al. (2024) 2025 ICLR 25
Moirai-MoE Liu et al. (2024b) 2024 ICML 25
VisionTS Chen et al. (2024) 2024 ICML 25
LightGTS Wang et al. (2025a) 2025 ICML 25
Sundial Liu et al. (2025) 2025 ICML 25
TimesFM-IC Faw et al. (2025) 2025 ICML 25
ROSE Wang et al. (2025b) 2025 ICML 25
FlowState Graf et al. (2025) 2025 NeurIPS 25 Workshop BERT2S
TiRex Auer et al. (2025) 2025 NeurIPS 25
Toto Cohen et al. (2025) 2025 NeurIPS 25
Pre-training datasets are used to fit a model to general time series patterns, such as, trends and seasonality. They are employed for initial training only and, hence, should be excluded from any evaluation activities. Train/test datasets, also sometime called in-domain or in-distribution datasets in the context of TSFMs, are datasets used to fine-tune a model to domain-specific patterns. They are split at a defined point in time into training and test sets and the model is fitted on the training data and evaluated on the test data (Hyndman and Athanasopoulos, 2021; Kapoor and Narayanan, 2023). Zero-shot datasets are used for evaluation of TSFMs on unseen data and should therefore be excluded from any training and fine-tuning. In a zero-shot setting, a model is given a short historical context of an unseen time series and it is assessed how well it can predict future values of this series.